A small rectangular park measures 200 m by 300 m. To the nearest meter, how much shorter is it if you walk along a diagonal to get to the opposite corner than if you walk along two adjacent sides?
Answer Choices
189
139
239
100

Answer by nerdybill(7384)

A small rectangular park measures 200 m by 300 m. To the nearest meter, how much shorter is it if you walk along a diagonal to get to the opposite corner than if you walk along two adjacent sides?
.
Length if you walk along two adjacent sides:
200 + 300 = 500m
.
Length of diagonal (applying Pythagorean theorem):
d^2 = 200^2 + 300^2
d^2 = 40000 + 90000
d^2 = 130000
d = 360.6
d = 361
.
500-361 = 139